The following company information is​ available:

Average accounts receivable $ 200,000

Average inventories 130,000

Average accounts payable 140,000

Net sales 10,000,000

Cost of goods sold 1,800,000

Purchases 1,140,000

What is the​ company's operating cycle​ (in days)?​ (Assume all sales are made on credit. Round your answer to the nearest​ day.)

A. 7 days

B. 26 days

C. 5 days

D. 34 days
Answer #1

Company's operating cycle = The days sales in inventory+average collection period)

Days Sales in inventory= 365 days/ Inventory turnover ratio

Inventory turnover ratio = Cost of goods sold/Average inventory.(1800000/130000=13.84)

:. Days sales in inventory = 365/13.84 = 26.37days..................(1)

Average collection period = 365/account receivable turnover ratio

Account receivable turnover ratio = Net credit sales/average account receivable (10000000/200000 = 50)

:. Average collection period=365/50 = 7.3days............(2)

Therefore company's operating cycle=(1)+(2) = (26.37+7.3) = 34 days.

So option D is correct answer.